OpenStudy (annjewell1998):

Solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair (a, b). a - b = -5 and a + 2b = 1 a (3, 2) b (3, -2) c (-3, -2) d (-3, 2) e no solution

OpenStudy (phi):

I would do a+2b=1 -(a-b = -5) i.e. subtract the lower equation from the top equation